OpenStudy (armyrangers):

Find the circumference and area of a circle with a diameter of 22 inches. Leave your answers in terms of pi. C = 11π; A = 44π C = 22π; A = 44π C = 11π; A = 121π C = 22π; A = 121π

OpenStudy (anonymous):

C= 22(pi); A=121(pi)

OpenStudy (texaschic101):

Circumference = pi * diameter....so circumference = pi * 22 Area = pi * r^2......keep in mind that r = 1/2d...so r = 11. Area = pi * 11^2.....Area = pi * 121

OpenStudy (armyrangers):

so is it c=11(pi) A=121(pi)

OpenStudy (texaschic101):

C = 22(pi) and A = 121(pi)

OpenStudy (armyrangers):

ok thanks

OpenStudy (texaschic101):

sure thing :)